Atsie (atsie):

Write the linear equation in standard form. 6 - y = 2x I just need to make sure I'm doing this right. :) If I add 2 to both sides it'd probably come out as 6 - y = 2y + 2y Then if I were to simplify it'd come out to... 2x + y = 6 Is this correct, or did I miss something? @phi

OpenStudy (math&ing001):

Standard form looks like this: y = mx + b It's like you're solving for y.

OpenStudy (sweetburger):

standard form is Ax+By = C y=mx+b is slope intercept form

Atsie (atsie):

I'm afraid the standard form I'm given in my chapter is \[Ax + By = C\] Not Y = mx + b

OpenStudy (math&ing001):

You're actually right !

OpenStudy (math&ing001):

My bad !

Atsie (atsie):

Who is? :P

OpenStudy (math&ing001):

@Atsie what you wrote is correct =)

OpenStudy (sweetburger):

@Atsie the answer you came to in your original question does appear to be correct.

Atsie (atsie):

Thank you both! You've been swell! @sweetburger and @math&ing001